# Fourier Boundary Features Network with Wider Catchers for Glass Segmentation

### v1

@@ III-D Fourier Convolution Controller
In the above, it is evident that adjacent inner sums are subtracted with an interval of 1 during the process of conducting FFT, signifying the subtraction of feature point values within the complex domain. Due to the images have naturally sparse representations [28], we only use the even symmetry components of each channel, the real part is reserved. Then, feature points exhibiting significant differences will retain larger positive or negative values within the feature map tensor. Naturally, the influence of these negative values will be mitigated by the ReLU function. For convenience, we use a one-dimensional convolutional layer $Conv$ to explain how FFT affects convolutional neural network learning.
Assuming the convolutional layer $Conv$ has three neurons which share the weights $W^{*}$, the feature extraction process of convolution can be expressed as:
$Input\cdot W^{*}=Output\cdot W^{*},$ (25)
where the $Input$ comes from the output of the previous layer, and the $Output$ is the features we want to extract. Specifically, what we want to obtain are boundary features, because the boundary features are crucial for glass segmentation. During network training, the correct weights $W^{*}$ can map the specific values of the $Input$ to more obvious values in $Output$. The loss function usually calculates the difference between the prediction value and the ground truth, but its essence is to calculate the errors between the ideal features and the extracted features by the network. We use $\mathscr{E}$ and $\mathscr{N}_{i}(i\in[0,1,2])$ to represent the errors and neural nodes, then we obtain the relationship between errors and weights:
$\frac{\delta\mathscr{E}}{\delta w_{i}}=\frac{\delta\mathscr{N}_{0}}{\delta w_{i}}\frac{\delta\mathscr{E}}{\delta\mathscr{N}_{0}}+\frac{\delta\mathscr{N}_{1}}{\delta w_{i}}\frac{\delta\mathscr{E}}{\delta\mathscr{N}_{1}}+\frac{\delta\mathscr{N}_{2}}{\delta w_{i}}\frac{\delta\mathscr{E}}{\delta\mathscr{N}_{2}},w_{i}\in W.$ (26)
The $\mathscr{E}$ depends on the tensor values difference between the $Output$ and the $Input$. If the tensor values of the $Input$ are close to the $Output$, the $\mathscr{E}$ will be more minor, then the fewer times it will take to update $W$. Since FFT extracts the edge information of the feature map, when the edge information is fused with the original feature map, the tensor values errors will be more minor. Due to convolution is a stretching of the original image, the original image $I^{*}$ also satisfies the above explanation, the Proposition 1 is proved.
@@ IV-A Experiment Setting
Implementation Details. The method outlined in our paper is implemented using the PyTorch framework on an NVIDIA RTX 4090 GPU equipped with 24GB memory. The three public glass segmentation datasets are optimized using an SGD optimizer with a weight decay of $5\times 10^{-4}$, an initial learning rate of 0.001, and a momentum of 0.9. Throughout the training process, we utilize the polynomial learning rate policy to gradually adjust the initial learning rate, denoted as $(1-\frac{{iter}}{{max_{iter}}})^{0.9}$.
@@ IV-B Comparison Results
In this study, to verify the effectiveness and robustness of the proposed method for the task of glass segmentation, the method is compared with a variety of methods in several related fields with few glass segmentation methods at present. We conduct extensive experiments to compare the segmentation performance of our work and state-of-the-art methods including PraNet [29], U-Net [30], and Transfuse [31] on medical image segmentation methods. DSS [32], PiCANet [33], RAS [34], CPD [35], EGNet [36], F3Net [37], MINet-R [38], ITSD [39], LDF [40] and U2-Net [41] on salient object detection methods. ICNet [42], PSPNet [43], DeepLabv3+ [44], DenseASPP [45], BiSeNet [46], DANet [47], CCNet [48], GFFNet [49], SFNet [50], FaPN [51] and PIDNet [52] on semantic segmentation methods. DSC [53], DSD [54], and BDRAR [55] on shadow detection methods. SINet [56], PFNet [57], and RankNet [58] on camouflaged object detection methods. MirrorNet [59] and PMD [1] on mirror segmentation methods. TransLab [5] and Trans2Seg [60] on transparent object segmentation methods. GDNet [9], GSD [2], EBLNet [6], PGSNet [14], and RFENet [61] on glass segmentation methods. To be fair, we retrain their open codes on the three datasets for effective comparison.
@@ IV-C Ablation Study
TABLE IV: Quantitative ablation results of CTA and FCC. The Base-Y means FBWC without CTA and FCC.
@@
A Base-Y $+$ FCC 83.48 0.068 6.73
Trans10K-stuff B Base-Y $+$ CTA 86.91 0.058 5.54
@@
TABLE V: Quantitative ablation results of FBWC without CTA on different number of CUs. The Base-Z means FBWC without WCC, CTA and FCC.
@@
Based on Table IV, we observe that the segmentation performance on all three datasets decreased without CTA to supplement the focused semantic features and maintain the integrity of semantic information in the edge constraint region. From Figure 9, it can be vividly seen that CTA can effectively filter the reflected strong light noise, fully focus the semantic information in the target region and reduce the dispersion of segmentation results. At the same time, we introduce the Fourier Transform on a variable branch to FCC, which is conducive to further strengthening the boundary information, increasing the boundary sensitivity of the model, and improving the segmentation accuracy. Visual results in Figure 10 show the importance of the FCC, which observes a sharp gradient decline of edge pixels in the conversion process of the frequency domain and spatial domain. Finally, the boundary information is integrated and strengthened to effectively divide the boundary between glass and reality.
@@
TABLE VI: Quantitative ablation results of FBWC without FCC on different number of CUs. The Base-Z means FBWC without WCC, CTA and FCC.
@@
TABLE VII: Quantitative ablation results of FBWC without CTA and FCC on different number of CUs. The Base-Z means FBWC without WCC, CTA and FCC.
@@
It can be observed from the data shown in Table V, under GDD dataset, the network segmentation performance of the FBWC without CTA experimental group becomes better with the increase of the number of CUs, and the best advantage occurs when the number of CUs is 4. When the number of CUs is 5, the network segmentation performance decreases to 86.39% IoU and 0.073 MAE. By horizontal contrast among Table V, Table VI, and Table VII, when the number of CUs is constant, the experimental group without both CTA and FCC in FBWC performs the most poorly compared with the other two groups. To be more specific, we observed that when the number of CUs was 3 in the HSO dataset, the network segmentation results without CTA and FCC are 70.69% IoU and 0.121 MAE, which are worse than the others with the same number of CUs. At the same time, under the experimental settings above mentioned, the segmentation performance results shown in Table V and Table VII are lower than those shown in Table VI. This reflects the importance and effectiveness of the FCC in the FBWC.
The experimental phenomena mentioned above is also reflected among three datasets with different numbers of CUs. When the number of CUs is 4, the overall network can have better segmentation performance. Besides, when the proposed method CTA and FCC in the FBWC are absent, the segmentation performance will decline.

### v2

@@ III-D Fourier Convolution Controller
According to Formula (22) and (23), there exists a recursive subtraction with a window size of 2 and a step size of 1 during the process of conducting FFT, which means that the difference between odd and even components is magnified in the frequency domain. Due to the images having naturally sparse representations [29], we only use the even symmetry components of each channel and the real part is reserved. According to [30], the IFFT is not required while the imaginary part is not used and the regular geometric structure feature in the pixel domain will be reserved. Then the boundary of glass in the pixel domain which is a regular geometric structure can be magnified in the frequency domain. However, using only frequency domain features may result in the loss of some refinement features in the pixel domain. To solve this problem, we fused the two heterogeneous source data in the frequency domain and spatial domain.
Proposition 1. Assuming $I^{*}$ is an image with regular geometric boundaries. For any continuous piecewise linear function subset $\Omega$ and any $f\in\Omega$, any convolutional neural network ( CNN ) $f_{J}^{\mathbf{w},\mathbf{b}}$ with depth $J\in\mathbb{Z^{+}}$ and ReLU function exists:
$\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(FFT(I^{*}).real+I^{*})\right\|_{\Omega}$ (24)
$<\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|_{\Omega}.$
Proof. For convenience, we consider the binary case, for simplicity, which is also suitable to the case in our study to explain how FFT affects CNN learning. Denote $D_{0}$ and $D_{1}$ to be the sets which contained the signals from two classes, and there is a positive gap between $D_{0}$ and $D_{1}$. That is, there exists some $d_{0}>0$ such as:
$d\left(D_{0},D_{1}\right)=\left\{\left\|x_{0}-x_{1}\right\|\mid x_{0}\in D_{0},x_{1}\in D_{1}\right\}=d_{0}>0.$ (25)
Assume that there is an ideal function $f$ such as:
$f(x)=0,\quad\mbox{ if }x\in D_{0}\quad\mbox{ and }\quad f(x)=1,\quad\mbox{ if }x\in D_{1}.$ (26)
For example, the $f(x)$ can be expressed as:
$f(x)=\begin{cases}con_{1}\cdot x+bias_{1},&x\in D_{0}\\ con_{2}\cdot x+bias_{2},&x\in D_{1},\end{cases}$ (27)
where $con_{1}$ and $con_{2}$ represent the coefficients, at the same time, the $bias_{1}$ and $bias_{2}$ represent the biases. The $f$, which is the operation of the convolutional layer, is composed of multiplication and addition. Then we can apply a CNN consisting of at least one convolution layer to approximate Formula (27). To increase the approximating ability, we usually introduce a nonlinear activation function. Since FFT in our case will produce some unwanted negative values, we use ReLU as the nonlinear function. Further, assuming $D_{0}\cup D_{1}=I^{*}$, the $I^{*}$ represents an image with regular geometric boundaries.
We input the image $I^{*}$ to a CNN with depth $J\in\mathbb{Z^{+}}$, and provided $S$ samples which are homologous to $I^{*}$ during the training. The approximating continuous piecewise linear function $f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}$ is obtained via solving the empirical minimization:
$\min_{\mathbf{w},\mathbf{b}}\frac{1}{S}\sum_{s=1}^{S}\left(f\left(I_{s}\right)-f_{J}^{\mathbf{w},\mathbf{b}}(I_{s})\right)^{2}.$ (28)
The approximating accuracy can be calculated as follows:
[…]
where $\mu$ is a probability measure on $I^{*}$. Direct estimation of Formula (29) is difficult due to the additional complication of sampling errors, thus most existing approaches estimate Formula (29) by separating it from the approximation error via triangle inequality:
$\left\|f(I^{*})-f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(I^{*})\right\|\leq\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|$ (30)
$+\left\|f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|,\forall h=f_{J}^{\mathbf{w},\mathbf{b}}\in\mathcal{H}^{J},$
where $\mathcal{H}^{J}$ represents the CNN space:
$\mathcal{H}^{J}=\left\{h=f_{J}^{\mathbf{w},\mathbf{b}}\mid\mathbf{w}\in\Phi_{\mathbf{w}}^{J},\mathbf{b}\in\Phi_{\mathbf{b}}^{J}\right\},$ (31)
where the $\Phi_{\mathbf{w}}^{J}=\cup_{\tau=1}^{J}\cup_{\upsilon=1}^{s_{\tau}}\left\{w^{\tau\upsilon}\right\}$ and $\Phi_{\mathbf{b}}^{J}=\cup_{\tau=1}^{J}\cup_{\upsilon=1}^{s_{\tau}}\left\{b^{\tau\upsilon}\right\}$ represent weight space and bias space, respectively. And the $s_{\tau}$ represents the convolution kernel size of $\tau$-th convolution layer. Since we consider that the CNN has converged, we no longer consider the value of $\left\|f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|$, which means $f_{J}^{\mathbf{w},\mathbf{b}}$ can be regarded as $f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}$ Then the approximation accuracy only needs to consider $\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|$.
According to Formula (25), if $f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}$ approximates to $f$, it should have the following properties:
$\|f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(x_{0})-f_{J}^{\mathbf{w},\mathbf{b}}(x_{1})\|\approx\|f(x_{0})-f(x_{1})\|\geq\|x_{0}-x_{1}\|.$ (32)
We use $Conv_{1}(\sigma(x;\Phi_{\mathbf{w}}^{1},\Phi_{\mathbf{b}}^{1}))$ to represent one convolution layer with ReLU function $\sigma$, then we can obtain:
$\|Conv_{1}(\sigma(x_{0};\Phi_{\mathbf{w^{*}}}^{1},\Phi_{\mathbf{b^{*}}}^{1}))-Conv_{1}(\sigma(x_{1};\Phi_{\mathbf{w^{*}}}^{1},\Phi_{\mathbf{b^{*}}}^{1}))\|$ (33)
$\geq\|x_{0}-x_{1}\|.$
For a CNN with two convolution layers, we have:
$\|Conv_{2}(\sigma(\textbf{O}_{1}^{x_{0}};\Phi_{\mathbf{w^{*}}}^{2},\Phi_{\mathbf{b^{*}}}^{2}))-Conv_{2}(\sigma(\textbf{O}_{1}^{x_{1}};\Phi_{\mathbf{w^{*}}}^{2},\Phi_{\mathbf{b^{*}}}^{2}))\|$ (34)
$\geq\|\textbf{O}_{1}^{x_{0}}-\textbf{O}_{1}^{x_{1}}\|.$
where $\textbf{O}_{1}^{x_{0}}$ and $\textbf{O}_{1}^{x_{1}}$ represent the output of first layer convolution with $x_{0}$ and $x_{1}$ as the original input, respectively. Then, for $f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}$, the input of the $J$-th layer is the output of the $(J-1)$-th layer, we can obtain the general form:
$\|Conv_{J}(\sigma(\textbf{O}_{J-1}^{x_{0}};\Phi_{\mathbf{w^{*}}}^{J},\Phi_{\mathbf{b^{*}}}^{J}))-Conv_{J}(\sigma(\textbf{O}_{J-1}^{x_{1}};\Phi_{\mathbf{w^{*}}}^{J},\Phi_{\mathbf{b^{*}}}^{J}))\|$ (35)
$\geq\|\textbf{O}_{J-1}^{x_{0}}-\textbf{O}_{J-1}^{x_{1}}\|\geq\cdots\geq\|\textbf{O}_{1}^{x_{0}}-\textbf{O}_{1}^{x_{1}}\|\geq\|x_{0}-x_{1}\|.$
If there exists $p\in{\{\|\textbf{O}_{J-1}^{x_{0}}-\textbf{O}_{J-1}^{x_{1}}\|,\dots,\|\textbf{O}_{1}^{x_{0}}-\textbf{O}_{1}^{x_{1}}\|\}}$ satisfying $\|\|f(x_{0})-f(x_{1})\|-p\|\leq\mathcal{E}$, where $\mathcal{E}$ means a minimum, we will consider $f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}$ approximates to $f$.
According to [31] and [32], we solved the Formula (28) by updating $\mathbf{w}$ and $\mathbf{b}$ through backpropagation errors, and the scope of each update is limited. After a finite number of iterations for Formula (28), the lower bound of $\|f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(x_{0})-f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(x_{1})\|$ is still determined by $\|x_{0}-x_{1}\|$, which is as shown in Formula (32). In our case, the positive gap is established by the border of glass. The pixel values within the boundry belong to $D_{1}$ and the pixel values outside the boundary belong to $D_{0}$. Then we replace $I^{*}$ by $(FFT(I^{*}).real+I^{*})$ as the input of $f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}$, the Formula (30) can be reformed as:
$\left\|f(I^{*})-f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(FFT(I^{*}).real+I^{*})\right\|$ (36)
$\leq\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(FFT(I^{*}).real+I^{*}))\right\|$
$+\left\|f_{J}^{\mathbf{w^{*}},\mathbf{b^{*}}}(FFT(I^{*}).real+I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(FFT(I^{*}).real+I^{*})\right\|$
$,\forall h=f_{J}^{\mathbf{w},\mathbf{b}}\in\mathcal{H}^{J}.$
In the above, the FFT can amplify the differences of the pixel around the boundary in frequency domain, then we can obtain:
$mean(\|x_{0}^{fft}-x_{1}^{fft}\|)>mean(\|x_{0}-x_{1}\|),$ (37)
$(x_{0}^{fft},x_{1}^{fft})\in(FFT(I^{*}).real+I^{*}),\quad(x_{0},x_{1})\in I^{*}.$
After $(FFT(I^{*}).real+I^{*})$ and $I^{*}$ are utilized as the inputs of the CNN, we used $p^{fft*}$ and $p^{*}$ to represent the maximum value in ${\{\|\textbf{O}_{J-1}^{x_{0}}-\textbf{O}_{J-1}^{x_{1}}\|,\dots,\|\textbf{O}_{1}^{x_{0}}-\textbf{O}_{1}^{x_{1}}\|\}}$, respectively. According to Formula (37), we can obtain:
$\|\|f(x_{0})-f(x_{1})\|-p^{fft^{*}}\|<\|\|f(x_{0})-f(x_{1})\|-p^{*}\|$ (38)
which means:
$\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(FFT(I^{*}).real+I^{*}))\right\|<\left\|f(I^{*})-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|.$ (39)
Then the Proposition 1 is proved.
@@ IV-A Experiment Setting
Implementation Details. The method outlined in our paper is implemented using the PyTorch framework on an NVIDIA RTX 4090 GPU equipped with 24GB memory. We run for 200 epochs with a batch size of 6 for the GPU. Before fine-tuning our models, we utilize pre-training weight as most of previous works, which is trained by ImageNet[33]. The three public glass segmentation datasets are optimized using an SGD optimizer with a weight decay of $5\times 10^{-4}$, an initial learning rate of 0.001, and a momentum of 0.9. Throughout the training process, we utilize the polynomial learning rate policy to gradually adjust the initial learning rate, denoted as $(1-\frac{{iter}}{{max_{iter}}})^{0.9}$.
@@ IV-B Comparison Results
In this study, to verify the effectiveness and robustness of the proposed method for the task of glass segmentation, the method is compared with a variety of methods in several related fields with few glass segmentation methods at present. We conduct extensive experiments to compare the segmentation performance of our work and state-of-the-art methods including PraNet [34], U-Net [35], and Transfuse [36] on medical image segmentation methods. DSS [37], PiCANet [38], RAS [39], CPD [40], EGNet [41], F3Net [42], MINet-R [43], ITSD [44], LDF [45], U2-Net [46] and Spider [47] on salient object detection methods. ICNet [48], PSPNet [49], DeepLabv3+ [50], DenseASPP [51], BiSeNet [52], DANet [53], CCNet [54], GFFNet [55], SFNet [56], FaPN [57], PIDNet [58] and BiRefNet [59] on semantic segmentation methods. DSC [60], DSD [61], and BDRAR [62] on shadow detection methods. SINet [63], PFNet [64], and RankNet [65] on camouflaged object detection methods. MirrorNet [66] and PMD [1] on mirror segmentation methods. TransLab [5] and Trans2Seg [67] on transparent object segmentation methods. GDNet [9], GSD [2], EBLNet [6], PGSNet [14], RFENet [68] and GateNet[69] on glass segmentation methods. To be fair, we retrained either their open codes or the implementations with recommended parameter settings on the three datasets for effective comparison.
@@ IV-C2 Trough Point
TABLE IV: Quantitative ablation results of BC Loss and AM Loss.
@@
A BC Loss (w/o) 83.08 0.068 6.75
Trans10K-stuff B AM Loss (w/o) 87.97 0.051 4.97
@@
TABLE V: Quantitative ablation results of CA, CTA, SCC and FCC. The Base-Y means FBWC without CA, CTA, SCC and FCC.
Datasets Methods Results
[…]
A Base-Y $+$ SCC 79.93 0.092 9.05
B Base-Y $+$ FCC 81.66 0.079 8.76
HSO C Base-Y $+$ CA 83.78 0.086 7.73
D Base-Y $+$ CTA 84.07 0.079 7.61
E FBWC 87.59 0.059 6.27
A Base-Y $+$ SCC 81.57 0.083 6.98
B Base-Y $+$ FCC 83.48 0.068 6.73
Trans10K-stuff C Base-Y $+$ CA 85.33 0.062 5.68
D Base-Y $+$ CTA 86.91 0.058 5.54
E FBWC 89.74 0.040 4.51
A Base-Y $+$ SCC 88.64 0.065 5.45
B Base-Y $+$ FCC 90.85 0.058 5.29
GDD C Base-Y $+$ CA 91.98 0.043 4.52
D Base-Y $+$ CTA 92.79 0.037 4.29
E FBWC 93.17 0.035 4.03
TABLE VI: Quantitative ablation results of FBWC without CTA on different number of CUs. The Base-Z means FBWC without WCC, CTA and FCC.
@@
IV-C3 Designed Module and Loss
From Table IV, we can observe the supervisory effects of BC Loss and AM Loss on the model. Specifically, when FBWC has no boundary constraints, the model’s ability to learn boundary features of the segmentation target weakens, resulting in a significant decrease in model performance. In addition, AM Loss can effectively improve the semantic feature consistency of large-scale segmentation regions in the model, thereby enhancing the coherence and naturalness of the segmentation results.
Based on Table V, we observe that the segmentation performance on all three datasets decreased without CTA to supplement the focused semantic features and maintain the integrity of semantic information in the edge constraint region. From Figure 9, it can be vividly seen that CTA can effectively filter the reflected strong light noise, fully focus the semantic information in the target region and reduce the dispersion of segmentation results. In order to verify the effectiveness of transposition operation in CTA, we removed the transposition process in the Cross Transpose Attention (CTA) module and made it a Cross Attention (CA) module. From Table V, it can be seen that the comprehensive performance of the CTA module on the three datasets is better than that of CA, indicating that transposition operation can establish effective global correlations among glass semantic features and boost segmentation performance.
At the same time, we introduce the Fourier Transform on a variable branch to FCC, which is conducive to further strengthening the boundary information, increasing the boundary sensitivity of the model, and improving the segmentation accuracy. Visual results in Figure 10 show the importance of the FCC, which observes a sharp gradient decline of edge pixels in the conversion process of the frequency domain and spatial domain. Finally, the boundary information is integrated and strengthened to effectively divide the boundary between glass and reality. Moreover, We designed relevant ablation experiments for this, specifically replacing FFT with a standard convolution operation to obtain Standard Convolution Controller (SCC). From Table V, it can be seen that SCC performs worse than FCC in the datasets HSO, Trans10K-stuff and GDD.
@@ IV-C3 Designed Module and Loss
TABLE VII: Quantitative ablation results of FBWC without FCC on different number of CUs. The Base-Z means FBWC without WCC, CTA and FCC.
@@
TABLE VIII: Quantitative ablation results of FBWC without CTA and FCC on different number of CUs. The Base-Z means FBWC without WCC, CTA and FCC.
@@
It can be observed from the data shown in Table VI, under GDD dataset, the network segmentation performance of the FBWC without CTA experimental group becomes better with the increase of the number of CUs, and the best advantage occurs when the number of CUs is 4. When the number of CUs is 5, the network segmentation performance decreases to 86.39% IoU and 0.073 MAE. By horizontal contrast among Table VI, Table VII, and Table VIII, when the number of CUs is constant, the experimental group without both CTA and FCC in FBWC performs the most poorly compared with the other two groups. To be more specific, we observed that when the number of CUs was 3 in the HSO dataset, the network segmentation results without CTA and FCC are 70.69% IoU and 0.121 MAE, which are worse than the others with the same number of CUs. At the same time, under the experimental settings above mentioned, the segmentation performance results shown in Table VI and Table VIII are lower than those shown in Table VII. This reflects the importance and effectiveness of the FCC in the FBWC.
The experimental phenomena mentioned above is also reflected among three datasets with different numbers of CUs. When the number of CUs is 4, the overall network can have better segmentation performance.
Figure 11: Visual results of FBWC on night scene .
Figure 12: Visual results of FBWC on noise condition.
Figure 13: Visual results of FBWC on occlusion condition.
IV-D Discussion and Analysis
IV-D1 Night Scenes
In night scenes, due to poor lighting conditions, the boundary information in the image becomes blurry, and the semantic features of large glass areas are not significant. To address this issue, the proposed FBWC could effectively segment glass regions as shown in Figure 11.
For the problem of boundary blurring under low light conditions, the WCC and continuous strong boundary constraints can effectively capture boundary features while preventing the dispersion of fine-grained boundary semantic information. Meanwhile, CTA enhances the consistency of glass regions through attention mechanisms. Finally, the boundary information and coarse-grained semantic features are further highlighted and effectively integrated through FCC. Based on the above content, FBWC can effectively cope with semantic segmentation tasks of glass images in night scenes. In addition, the datasets utilized in this work contain real-world glass images under low light conditions, providing sample diversity and further enhancing the model’s generalization ability.
IV-D2 Non-ideal Conditions
Similarly, when dealing with non-ideal conditions such as noise and occlusion in the glass semantic segmentation task, boundary and glass region features are easily affected by noise interference and are difficult to extract. The segmentation results of the proposed FBWC under Gaussian noise, Salt&Pepper noise, Poisson noise, ColorJitter, and Foggy noise are shown in Figure 12, and the segmentation results under occlusion conditions are shown in Figure 13.
In order to effectively address noise and occlusion issues, FBWC applies WCC to avoid the over-capturing of large glass areas, which is beneficial for maintaining the integrity of glass areas under occlusion, light spots, and noise interference. CTA is further designed to achieve effective segmentation of non-occluded areas. Simultaneously, through multi-level boundary loss constraints and effective boundary enhancement of FFT, the boundary information in shallow structures has been significantly optimized, further improving the anti-interference ability and robustness of FBWC.
IV-D3 Complexity and Efficiency
We have indicated the parameter count for each module in Figure 2. It is worth mentioning that the parameters of the whole FBWC are 48.7M and the backbone WCC we designed is 38.38M. Regarding the backbone, WCC reduces the parameters of 47.62M, 99.62M and 5.62M compared to ViT-Base[70], VGG16[71] and ResNeXt101[72], whose are used in the popular methods of glass segmentation. In addition, the parameters of the proposed FCC and SegHead are 2.37M, which does not significantly increase the parameters of the FBWC. This optimization not only reduces the computational complexity of the model but also improves the training and inference efficiency of the model. By reducing the number of parameters, WCC reduces the storage and computational burden of the model while maintaining high performance.